8. A spring stretches 3 cm when a 5 g object is hung from it. The object is replaced with a block of mass 50 g which oscillates in simple harmonic motion. What is the period of the motion? 9. It is known that a load with a mass of 25 g will stretch a spring 5 cm. The spring is stretched an additional 30 cm, and then released. (a) What is the springs constant? (b) What is the period and frequency of the oscillations? 10. How would you determine the mass of an object on the space station in space?
Answer #1

here,

8)

let the spring constant be K

when m1 = 5 g = 0.005 kg hangs from the spring

x = 3 cm = 0.03 m

equating the forces vertically

K * x = m * g

K * 0.03 = 0.005 * 9.81

solving for K

K = 1.635 N/m

New mass attached , m2 = 50 g = 0.05 kg

the period of motion , T = 2*pi*sqrt(m/K)

T = 2*pi*sqrt(0.05 /1.635)

T = 1.1 s